the mathematics teacher claims that the mean iq of statistics students is 110 with standard deviation of 12. The mean IQ of 28 randomly selected statistics students is 112 . Test the difference of the population and sample means at 5%level of significance.

1
Expert's answer
2022-06-02T17:00:10-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=110"

"H_1:\mu\not=110"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z:|z|>1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{112-110}{12/\sqrt{28}}=0.8919"

Since it is observed that "|z|=0.8919<1.96=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=2P(z>0.8919)=0.372447," and since "p= 0.372447>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 110, at the "\alpha = 0.05" significance level.